Math 2418 Linear Algebra Quiz #2

[10] 1.) Prove that matrix multiplication for SQUARE matrices is not commutative (I.e., give two specific square matrices with real numbers and show with these matrices AB ≠ BA).

[10] 2.) Given the following augmented matrix, solve

[ \begin{matrix}1 & 0 & 0 & 0 & 0 & −3 \\ 0 & 1 & 2 & 0 & 1 & 5 \\ 0 & 0 & 0 & 1 & −3 & 2 \end{matrix} ]